find the discriminant of y=x2-3x-5
Found 2 solutions by stanbon, mathstutor458:
Answer by stanbon(75887)

find the discriminant of y=x2-3x-5
---
b^2-4ac = 9-4*-5 = 29
====================
Cheers,
Stan H.
===============

Answer by mathstutor458(57)

given, y=x2-3x-5
discriminant formula = b2-4ac
here a=1(coefficient of x2)
b=-3(coefficient of x)
c=-5(constant)
discriminant = (-3)2-4*1*(-5)=9+20=29 --answer